OpenStudy (anonymous):

What does it mean to refer to a 20 in TV set or a 25 in TV set? Such units refer to the diagonal of the screen. A 35 in TV set also has a width of 28 inches. What is its height?

OpenStudy (anonymous):

Use that Pythagorean theorem \[28^{2}+x^{2}=35^{2}\]\[x^{2}=441\]\[x=21\]

OpenStudy (anonymous):

using Pythagorean theorem, \[\sqrt{(c^2 - b^2)} = a\] where a is your height, using the basic theorem

OpenStudy (anonymous):

Let the height be "x": Equation: x^2 + 28^2 = 35^2 x^2 = 35^2-28^2 x^2 = 441 x = 21 inches =================